A. At a certain temperature, the pH of a neutral solution is 7.67. What is the value of Kw at that temperature?

Express your answer numerically using two significant figures.

B.

What is the pH of a neutral solution at a temperature where Kw=9.9×10−14?

Express your answer numerically using two decimal places.
Answer #1

A)

use:
pH = -log [H+]
7.67 = -log [H+]
[H+] = 2.138*10^-8 M

Since solution is neutral,
[OH-] = [H+] = 2.138*10^-8 M

Use:
Kw = [H+][OH-]
= (2.138*10^-8)*(2.138*10^-8)
= 4.57*10^-16
Answer: 4.6*10^-16

B)
Kw = [H+][OH-]
Since solution is neutral, [H+] = [OH-]
So,
Kw = [H+]^2
9.9*10^-14 = [H+]^2
[H+] = 3.15*10^-7 M


use:
pH = -log [H+]
= -log (3.15*10^-7)
= 6.5017
Answer: 6.50